OpenStudy (anonymous):

Eliminate the “z” variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 Answer 6x – 5y = –13 4x + y = –14 4x – 3y = –14 6x + 5y = 13 .

OpenStudy (anonymous):

2x-y+2z=-15 2(2x-2y-z)=2(1) 2x-y+2z=-15 4x-4y-2z=2 add together 6x-5y=13

OpenStudy (anonymous):

6x-5y=-13*

hero (hero):

\[\color {green} {\huge{\checkmark}} \]